OpenStudy (anonymous):

Solve the differential equation\[\frac{dy}{dx} = 4 x^2 y^2\]with the condition that \(y( 2) = 4\)

OpenStudy (turingtest):

please do not close and repost the same question

OpenStudy (anonymous):

@fazeer , you are forgetting \ [ \] (without the space :-) )

OpenStudy (anonymous):

i am getting 4x^3y^2/3 + c

OpenStudy (anonymous):

it bumped me out

OpenStudy (turingtest):

bumped you out? only you or a moderator can close your Q, and I didn't do it o-0

OpenStudy (turingtest):

anyway, did I fix your post correctly?

OpenStudy (turingtest):

...

OpenStudy (anonymous):

what is the answer plz

OpenStudy (turingtest):

did I type it right?

OpenStudy (anonymous):

????

OpenStudy (turingtest):

is this a language problem we are having here? I'm trying to make sure I typed your problem right if I did the equation is separable and all you need to do is get like terms together and integrate

OpenStudy (turingtest):

...then apple the initial conditions to find the value of the integration constant C

OpenStudy (anonymous):

can you give me the answer, its not 4x^3y^2/3

OpenStudy (anonymous):

\[ \frac{dy}{dx} = 4x^2y^2 \implies \frac{dy}{y^2} = 4x^2 dx\]Integrate both sides. Your answer is wrong.

OpenStudy (turingtest):

and no we cannot just give you the answer, that is against the code of conduct of this site. I'm pretty surprised to see a differential equation student who's not familiar with wolfram though

OpenStudy (turingtest):

though I guess they give you these problems in basic calc sometimes....

OpenStudy (anonymous):

Can you integrate dy/ y^2 ?